OpenStudy (anonymous):

Choose the equation of the line that has a slope of -5/6 and contains the point (-6,4) A. y=-5/6 - 1 B. y= -5/6 + 1 C. y=5/6 - 1

OpenStudy (nurali):

y=mx+c (x,y)=(-6,4) m=-5/6 4=(-5/6)*(-6)+b can you solve for b now?